OpenStudy (anonymous):

Solve the equation. Check both solutions and only write the real solution(s). square root of 2x+3=x

OpenStudy (anonymous):

Is this what you mean \[\sqrt{2x + 3} = x \]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

We are going to solve it together. First square both sides that gives 2x + 3 = x^2

OpenStudy (anonymous):

does that make sense

OpenStudy (anonymous):

If it is, square both sides to get:\[x^{2} = 2x + 3 \rightarrow x^{2} - 2x - 3 = 0\]Remember that\[\Delta=b^{2} - 4ac\]In specific, we get:\[\Delta = 4 - (4*(-3)*1) = 16\]So the binomial has distinct real roots.

OpenStudy (anonymous):

@denyekwe what do i do after that?

OpenStudy (anonymous):

next you will want to subtract(2x + 3) from both sides of the equation.

OpenStudy (anonymous):

I chose (2x + 3), because i wanted to arrive to x^2 - 2x - 3 = 0

OpenStudy (anonymous):

then do i solve after this ?

OpenStudy (anonymous):

Yes, we are left with a quadratic equation.

OpenStudy (anonymous):

do i end up with x^2-2x=0?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

What did you do?

OpenStudy (anonymous):

x^2-2x=0 then added 3 to both sides

OpenStudy (anonymous):

If you add three to both sides we will have something like this. x^2 - 2x - 3 + 3 = 0 + 3

OpenStudy (anonymous):

Evaluating that will give us, x^2 - 2x = 3

OpenStudy (anonymous):

i'm sorry, that is what i got

OpenStudy (anonymous):

Thats fine. Everyone makes mistakes.

OpenStudy (anonymous):

To solve x^2 - 2x - 3 = 0 we have to treat it as a quadratic equation.

OpenStudy (anonymous):

Are you familiar with quadratic equations

OpenStudy (anonymous):

kinda

OpenStudy (anonymous):

I would love to teach you somethings about quadratic equations but i have something else i am working on. This link " http://www.khanacademy.org/ " has lots of FREE videos on solving quadratic equations. I suggest you visit it and watch the lectures.

OpenStudy (anonymous):

The list of videos is really long so i suggest that you do a find on the web page for the word "Algebra: Quadratics". This will take you to the section where videos for solving quadratic equations are located.

OpenStudy (anonymous):

But to finish up the problem we had, i will use the quadratic formular.

OpenStudy (anonymous):

(1)x^2 - (2)x - (3) = 0 a b c The quadratic formula \[x = (-b \pm \sqrt{b ^{2} - 4ac})\div2a\] from our equation, a = 1 b = -2 c = -3 \[x = (-(-2) \pm \sqrt{(-2) ^{2} - 4(1)(-3)})\div2(1)\] \[x = ( 2\pm \sqrt{4 + 12})\div2\] \[x = ( 2\pm \sqrt{16})\div2\] \[x = ( 2\pm 4)\div2\] this leaves us with x = (2-4)/2 = -1 x = (2+4)/2 = 3

OpenStudy (anonymous):

Going back to the original question. (2x+3)^(0.5) = x when we sub our x from above, we get (2(-1)+3)^(0.5) = (-1) (-2+3)^(0.5) = (-1) (1)^(0.5) = (-1) 1 = -1 ### Since this statement is not true x = -1, does not belong to my solution. (2(3)+3)^(0.5) = (3) (6+3)^(0.5) = (3) (9)^(0.5) = (3) 3 = 3 ### Since this statement is true x = 3, does belong to my solution.

OpenStudy (anonymous):

There is a special trick to do this problem. You can show that if a-b+c =0, then one root is -1 and the other is -c/a On a related matter, if a+b+c=0 then one root is 1 and the other is c/a. Both statement comes from the fact that the product of the two roots is c/a

OpenStudy (anonymous):

You can practice problems of roots of polynomials on http://www.saab.org/mathdrills/act.html

jhonyy9 (jhonyy9):

sqrt(2x+3)=x both sides squared 2x+3=x^2 x^2 -2x -3=0 (x-3)(x+1)=0 x-3=0 x=3 x+1=0 x=-1